OpenStudy (anonymous):

How do you solve this equation using substitution: 3x-y=14 2x+y=16

OpenStudy (anonymous):

the Y values cancel each other out so it becomes 5x=30, 5 divides into 30 6 times.

OpenStudy (anonymous):

using substitution, not elimination

OpenStudy (anonymous):

Or solve one of the equations for y, then plug it in to the other equation.

OpenStudy (anonymous):

Like y=-2x+16, then plug this into the other one: 3x-(-2x=16)=14 Solve for x, then place this the number back into the first equation.

OpenStudy (anonymous):

So 3x+2x-16=14 5x-16=14 x=6

OpenStudy (anonymous):

So x=6 and now place that back into the other equation: 2(6)+y=16 12+y=16 y=4